# Are Language Models More Like Libraries or Like Librarians? Bibliotechnism, the Novel Reference Problem, and the Attitudes of LLMs

### 10 Jan 2024 v1

@@ Abstract
Are LLMs cultural technologies like photocopiers or printing presses, which transmit information but cannot create new content? A challenge for this idea, which we call bibliotechnism, is that LLMs often do generate entirely novel text. We begin by defending bibliotechnism against this challenge, showing how novel text may be meaningful only in a derivative sense, so that the content of this generated text depends in an important sense on the content of original human text. We go on to present a different, novel challenge for bibliotechnism, stemming from examples in which LLMs generate “novel reference”, using novel names to refer to novel entities. Such examples could be smoothly explained if LLMs were not cultural technologies but possessed a limited form of agency (beliefs, desires, and intentions). According to interpretationism in the philosophy of mind, a system has beliefs, desires and intentions if and only if its behavior is well-explained by the hypothesis that it has such states. In line with this view, we argue that cases of novel reference provide evidence that LLMs do in fact have beliefs, desires, and intentions, and thus have a limited form of agency.
@@ 1 Introduction
Do modern LLMs have beliefs, desires, and intentions? Over the last few years, this question has received an enormous amount of attention (e.g., Hase et al., 2023; Mahowald et al., 2023; Bender and Koller, 2020; Shanahan et al., 2023; Bubeck et al., 2023). The hypothesis that LLMs do have these states is attractive in part because it offers a natural tool for explaining their behavior. It is standard to explain the complex behavior of humans and non-human animals in terms of what they think (believe), what they want (desire), and what they intend. If modern LLMs have beliefs, desires, and intentions, that is, if they are agents, we can employ the same explanations of their behavior.
@@
A challenge for those who deny that current LLMs are agents in this sense is to provide an alternative, equally powerful, explanation of their behavior. The psychologist Alison Gopnik and her coauthors have articulated a striking idea in this direction (Gopnik, 2022b, a; Yiu et al., 2023). In Gopnik’s view, LLMs are a “cultural technology”, like a library or a printing press. The writer Ted Chiang also gives voice to an idea in this vein: “Prompting it [the LLM] with text is something like searching over a library’s contents for passages that are close to the prompt, and sampling from what follows.” (Chiang, 2023). Cosma Shalizi, who has developed this idea in more technical detail (Shalizi, 2023), has dubbed the view “Gopnikism”. Because we will develop it in our own direction, we call it “bibliotechnism”, combining the Greek for “book” with the Greek for “skill”. According to bibliotechnism, LLMs are not agents; they are “just” cultural technologies, like books and libraries, for processing and querying written text.
@@
Can this view provide an explanation of the agent-like behavior of LLMs, which is sufficiently powerful to compete with the hypothesis that they have beliefs, desires and intentions? We address this question in a specific application, by examining the meaning-relevant behavior of LLMs (joining a growing body of work at the intersection of philosophy, cognitive science, and NLP; Bender and Koller, 2020; Andreas, 2022; Coelho Mollo and Millière, 2023; Chalmers, 2023; Mandelkern and Linzen, 2023; Piantadosi and Hill, 2022). We argue that if LLMs are “just” a cultural technology (and not agents in their own right) then the fact that their outputs refer to certain objects must in an important sense depend on the fact that their inputs refer to those objects. If LLMs’ reference were not of this “derivative” kind, then there would be an important sense in which they were not simply transmitting existing cultural knowledge, but generating new instances of reference, and perhaps even new claims.
@@
In normal cases, text produced by photocopiers and printing presses clearly has only derivative meaning, since it is simply a reproduction of human-generated input. But LLMs often produce apparently meaningful text, which is entirely novel. At first sight, this fact presents a serious challenge for bibliotechnism: if LLMs’ outputs can only be meaningful by piggybacking on human-generated originals, how could novel text that they generate be meaningful?
@@
We begin by responding to this challenge for bibliotechnism. Using n-grams as a toy model, and working up to more complex modern LLMs, we show how even entirely novel sentences produced by LLMs may nevertheless derive their meaningfulness from the meaningfulness of their inputs and thus be “derivatively meaningful”.
@@
We see this as a big step forward for bibliotechnism. But challenges for the proposal remain. Modern LLMs are not just capable of producing new sentences, they can also generate novel reference, using newly invented names apparently to refer to newly created objects. These new names cannot derive their reference from original text, since the name and its object are not associated in the data (if they occur there at all). We argue that responding to this Novel Reference Problem requires complicating bibliotechnism to such an extent that it calls into question the motivation for doing so.
@@
In particular, bibliotechnism’s answer to the Novel Reference Problem is a worse explanation of LLM behavior than the hypothesis that they have beliefs, desires, and intentions. According to many views in the philosophy of mind, this fact provides evidence that LLMs do have beliefs, desires, and intentions. This is most obviously true given interpretationism, according to which a system has beliefs, desires and intentions if and only if its behavior is best explained by the hypothesis that it is rational and that it has such states (e.g., Dennett, 1971; Davidson, 1973, 1986; Dennett, 1989; McCarthy, 1979). But the point holds for other prominent philosophical views as well. So, the problem of novel reference provides evidence that LLMs do have beliefs, desires and intentions, and thus are agents of a sort.
2 Prior Work on Meaning in LLMs
@@
A prominent line of argument has suggested that LLMs cannot produce reference without being “grounded” (e.g., Lake and Murphy, 2023; Bisk et al., 2020). Perhaps most influentially, Bender and Koller (2020) examine the question of whether LLMs use language meaningfully. They define “meaning” as a relation between expressions and communicative intents. They argue that LLMs cannot produce meaningful expressions because they cannot have communicative intents concerning objects they have not had perceptual contact with.
@@
Mandelkern and Linzen (2023) observe important connections between this debate and semantic externalism, a view of meaning which has been dominant in the philosophy of language since the 1980s. On a standard view (Kripke, 1980; Putnam, 1975; Burge, 1986), people can refer to Shakespeare without having been directly in touch with Shakespeare, by belonging to a community whose overall use of this word stands in an appropriate causal relationship to the poet. Mandelkern and Linzen accordingly argue that whether LLMs can refer to Shakespeare comes down to whether LLMs “belong to our speech community” (cf. Ostertag, 2023).
@@
These authors do not consider how questions about the meaningfulness of LLM-generated text relate to bibliotechnism. They also do not address how, if at all, simpler models like n-gram models can produce meaningful text. In fact, to the extent that they deliver verdicts about n-grams, their theories do not clearly imply that n-grams can produce meaningful text. By contrast, we will argue that n-gram models can clearly produce meaningful words. We will build on this account in the case of n-grams to offer a novel extension of bibliotechnism, showing how the view can accommodate the meaningfulness of entirely novel text generated by LLMs. We then introduce the “Novel Reference Problem”, a new challenge for bibliotechnism.
3 Background
@@
Three points about philosophical terminology will be important. Philosophers often distinguish between “reference” and “meaning”. As we will use the terms, any expression that refers has a meaning, although many meaningful expressions do not refer. For simplicity, the only words that we take to refer are meaningful common and proper nouns. It is uncontroversial that a normal use of the word “Shakespeare” refers to Shakespeare (and is meaningful). By contrast, in our (stipulative) usage, normal uses of the expression “was born” is meaningful, but it does not refer.
Second, there is a difference between the word “Shakespeare” and particular inscriptions or tokens of this word. If the word “Shakespeare” is written on a blackboard five times, there are five inscriptions of this one word on the blackboard. We assume inscriptions of words can refer and be meaningful.
Third, and finally, we will distinguish between “referring” that is done by an agent (“In his indirect way, Marlowe was referring to Queen Elizabeth.”), and referring that is done by particular inscriptions of words. Since our goal is to explore a view on which LLMs are not agents, we will be investigating the question of whether they can produce inscriptions which refer and are meaningful. We will not be assuming that they themselves can refer. [^1] Mandelkern and Linzen (2023) often move without comment between the question of whether a particular agent refers, and the question of whether particular inscriptions of words refer. But this distinction seems to us of key importance here, since it may take beliefs and desires to refer as an agent, but it does not take such attitudes to produce inscriptions which refer (as we will argue in a moment, photocopiers can do so, as can n-grams).
4 From Cultural Technology to Derivative Reference and Meaning
@@
Gopnik and other bibliotechnists understand cultural technologies, like books and libraries, as tools for the transmission and dissemination of information, allowing the accumulation of knowledge over large stretches of space, and, most notably, time. These technologies are, crucially, not themselves responsible for new ideas or information.
@@
We will say that the original inscriptions, which were created by the author immediately, are instances of basic reference and meaning, while the other inscriptions are instances of derivative reference and meaning. We stipulate, as part of the definition of these terms, that it is only agents (which we understand to mean entities with beliefs, desires, and intentions) who produce inscriptions which refer or are meaningful basically. Beyond this stipulation, the distinction between basic and derivative reference and meaning is rough, but we will only deal with clear examples of each category in what follows.
According to bibliotechnism, LLMs are not agents. So, according to bibliotechnism, they can only produce inscriptions which refer or are meaningful derivatively.
@@
5 Causal Connection and Derivative Reference: The Case of N-grams
In this section, we argue that derivative reference and meaning can be achieved by an appropriate causal connection between PrimaryData and GeneratedText, and show how this vindicates the idea that n-grams can produce derivatively meaningful words.
@@ 5 Causal Connection and Derivative Reference: The Case of N-grams
This observation already shows that large-n n-gram models (with $n$ sufficiently large, e.g. 1000, that they copy long stretches of their input) can produce meaningful inscriptions. Since such an n-gram model copies their input, their output can be meaningful, just as the photocopier’s is.
Matters are less straightforward for unigram models, which sample from a distribution of single words. We can think of this model as implemented by taking all of its PrimaryData, choosing word-inscriptions from the PrimaryData at random, and then copying the chosen inscription. The inscriptions the n-gram then produces are again just like those of a copier (or of the large-n model): they have a direct causal connection to the original inscriptions. As a result, if the model produces an inscription of “Shakespeare”, this inscription will be meaningful (and refer) derivatively, piggybacking on the meaning and reference of the original inscription of this word.
In this case, however, there is a new phenomenon, not exhibited in the case of the photocopier or large-n n-gram. Each of the inscriptions of the individual words produced by the unigram will be meaningful (and possibly refer), but it does not seem that inscriptions of complex expressions formed from these words will be. The vast majority of the time, the string of words the model produces will be gibberish, and uncontroversially meaningless. At low odds the unigram model will produce a “reasonable” string like “Shakespeare was born in 1564”. With even lower odds, such a reasonable string will be produced as an exact copy of an original string. But even in these latter cases, the fact that a meaningful string (or even a copy) is produced is a fluke, a complete accident. Since the model produces “appropriate” expressions in such a chaotic way, we judge that, even when the model produces an inscription of a string that would be meaningful if produced by a human in a normal way, this inscription does not have a meaning. The generated text is not appropriately causally connected to the PrimaryData to inherit “glue” that binds the complex expression together. One way to put this point would be to say that, while the inscription it produces may look like an inscription of a sentence (and a human who sees it may be able to conjure up a meaning associated with it), it is not really an inscription of a sentence, but just an inscription of words which could (in different circumstances) have made up a sentence. These inscriptions are like words blown together by the wind, or sand dunes blown into the shape of a sentence—or like Mandelkern and Linzen’s “ants formicating meaninglessly in the sand”.
6 LLMs do Produce Derivatively Meaningful Complex Expressions
@@
We will argue that it can, by arguing that there is another route to producing derivatively meaningful inscriptions of complex expressions, where the causal chain which traces from the inscription of individual words back to the original inscriptions is distinct from the causal chain which is responsible for the expression-level features which make a whole complex expression meaningful.
@@
To see the basic idea, suppose we have a rudimentary model, which, when fed some text, finds any inscriptions of names in this sentence (searching on the basis of a database) and then replaces each of these names uniformly with a name drawn at random from the distribution of all names in its PrimaryData. It seems plausible that in this case, the model would produce not just individual words which are derivatively meaningful, but a new sentence which would, as a whole, be derivatively meaningful. For instance, if we gave this model our Shakespeare sentence, and it produced “Barack Obama was born in 1564” this inscription would be false, but meaningful. Here, the causal history of the context and the causal history of the individual name printed are different, but the whole expression would still be derivatively meaningful. Plausibly, this is because the operation as a whole is causally sensitive to sentence-level features, in such a way as to reliably produce an intelligible sentence. This could be so, even if the resulting sentence has never been seen before.
This example shows that even novel text can be derivatively meaningful. It also suggests that, if LLMs are causally sensitive to high-level features of their PrimaryData, in such a way as to transmit those features to their GeneratedText, it is possible that even their entirely novel GeneratedText could be derivatively meaningful. We suggest that one such high-level feature is intelligibility. As we will understand the notion, intelligibility of expressions requires they they be at least quasi-grammatical (sentences with minor grammatical errors are often perfectly intelligible), but it requires more than just grammaticality, since not all grammatical sentences are meaningful (and hence not intelligible). If LLMs are causally sensitive to intelligibility, so that they produce intelligible outputs from intelligible inputs, then intelligible complex expressions in their GeneratedText will be derivatively meaningful, with individual words inheriting the meanings of the original inscriptions from which they were “copied”, and whole expressions inheriting the “glue” of intelligibility from the PrimaryData.
There is strong evidence that modern LLMs are in fact appropriately causally sensitive to the intelligibility of their PrimaryData. First, they overwhelmingly produce text which is intelligible to human users. This does not prove that they are causally sensitive to this property, but it is strong evidence that they are. Second, it seems extremely plausible that if LLMs were trained on gibberish, they would output gibberish. These two claims at least point toward the verdict that they causally transmit the intelligibility of their data.
@@
Intelligibility in our sense does not require truth or even sufficiently reliable production of the truth. False sentences like “Shakespeare was born in 2023” are intelligible. This is important because even the best LLMs at the time of writing are known to confabulate or fabricate information. But getting a fact wrong (e.g., saying Shakespeare was born in 2023 instead of 1564) is importantly different than if the model produces incoherent responses. If an LLM reliably responded to queries about Shakespeare’s birth with gibberish, this would at least be some evidence that it is not in fact causally sensitive to the intelligibility of its data in such a way as to generate derivatively meaningful complex expressions.
It is instructive to compare LLM-generated text to text generated by bigram or trigram models. Bigram and trigram models are trickier cases than unigrams because they do copy short complex phrases and might be statistically likely to combine them in ways that are more plausibly meaningful as a whole. For instance, a bigram model that outputs “Shakespeare wrote plays” in a way “knows” that “wrote” is a likely continuation for “Shakespeare” and that “plays” is a likely continuation for ”wrote”. But, besides the one mediated by the verb “wrote”, there is no causal connection between “Shakespeare” and “plays”. So, even when the model produces strings of sentence-length that are grammatical, and even when individual phrases may be judged meaningful, it seems that, as with unigram models, longer sentences should probably not be understood as meaningful (although this is more of a borderline case): they lack the straightforward “copy property” of higher-n n-gram models but also are not causally sensitive to intelligibility as modern LLMs are.
@@
Before closing this discussion, we want to offer one important clarification about the basis of our judgment that unigrams do not produce derivatively meaningful complex expressions. The basis for this judgment is not the fact that n-grams are only trained on words. It is instead because this training mechanism does not lead to causal sensitivity to relevant high-level features of their PrimaryData. To put this another way: we are not interested in a narrow form of “input-sensitivity”, but instead in a broader notion of causal sensitivity.
@@
This contrast can be illustrated by considering again a photocopier. The fact that a photocopier responds to (say) one or another aspect of the ink used to write original letters is irrelevant to the question of whether the tokens it produces refer—as long as this underlying low-level mechanism reliably produces inscriptions of actual words, that is, as long as the low-level mechanism leads to causal sensitivity to the right high-level features.

### 15 Feb 2024 v2

@@ Abstract
Are LLMs cultural technologies like photocopiers or printing presses, which transmit information but cannot create new content? A challenge for this idea, which we call bibliotechnism, is that LLMs often generate entirely novel text. We begin (Part I) with a sustained defense of bibliotechnism against this challenge showing how even entirely novel text may be meaningful only in a derivative sense, and arguing that, in particular, much novel text generated by LLMs is only derivatively meaningful. But we argue (Part II) that bibliotechnism faces a different, novel challenge, stemming from examples in which LLMs generate “novel reference”, using novel names to refer to novel entities. Such examples could be smoothly explained if LLMs were not cultural technologies but possessed a limited form of agency (beliefs, desires, and intentions). According to interpretationism in the philosophy of mind, a system has beliefs, desires and intentions if and only if its behavior is well explained by the hypothesis that it has such states. So, according to interpretationism, cases of novel reference provide evidence that LLMs have beliefs, desires, and intentions. Given that interpretationism is a live hypothesis about the nature of these states, we suggest that cases of novel reference provide evidence that LLMs do have beliefs, desires, and intentions.
@@ 1 Introduction
Do modern LLMs have beliefs, desires, and intentions? Over the last few years, this question has received an enormous amount of attention (e.g., Hase et al., 2023; Mahowald et al., 2023; Shanahan et al., 2023; Bubeck et al., 2023; Levinstein and Herrmann, 2023; Yildirim and Paul, 2023). The hypothesis that LLMs do have these states is attractive in part because it offers a natural tool for explaining their behavior. It is standard to explain the complex behavior of humans and non-human animals in terms of what they think (believe), what they want (desire), and what they intend. If modern LLMs have beliefs, desires, and intentions, we can employ the same explanations of their behavior.
@@
A challenge for those who deny that current LLMs have beliefs, desires, and intentions, is to provide an alternative, equally powerful, explanation of their behavior. The psychologist Alison Gopnik and her coauthors have articulated a striking idea in this direction (Gopnik, 2022b, a; Yiu et al., 2023). In Gopnik’s view, LLMs are a “cultural technology”, like a library or a printing press. The writer Ted Chiang also gives voice to an idea in this vein: “Prompting it [the LLM] with text is something like searching over a library’s contents for passages that are close to the prompt, and sampling from what follows.” (Chiang, 2023). Cosma Shalizi, who has developed this idea in more technical detail (Shalizi, 2023), has dubbed the view “Gopnikism”. Because we will develop it in our own direction, we call it “bibliotechnism”, combining the Greek for “book” with the Greek for “skill”. According to bibliotechnism, LLMs are not agents; they are “just” cultural technologies, like books and libraries, for processing and querying written text.
@@
Can this view provide an explanation of the agent-like behavior of LLMs, which is sufficiently powerful to compete with the hypothesis that they have beliefs, desires and intentions? We address this question in a specific application, by examining the meaning-relevant behavior of LLMs, joining a growing body of work at the intersection of philosophy, cognitive science, and NLP (Bender and Koller, 2020; Andreas, 2022; Coelho Mollo and Millière, 2023; Chalmers, 2023; Mandelkern and Linzen, 2023; Piantadosi and Hill, 2022; Millière and Buckner, 2024; Titus, 2024). We argue that if LLMs are “just” a cultural technology (and not agents in their own right) then the fact that their outputs refer to certain objects must in an important sense depend on the fact that their inputs refer to those objects. If LLMs’ reference were not of this “derivative” kind, then there would be an important sense in which they do not simply transmit existing cultural knowledge, but generate new instances of reference, and perhaps even new claims.
@@
In normal cases, text produced by photocopiers and printing presses clearly has only derivative meaning, since it is simply a reproduction of human-generated input. But LLMs often produce entirely novel text, which is still apparently meaningful. At first sight, this fact presents a serious challenge for bibliotechnism: if LLM-generated text can only be meaningful if it piggybacks on human-generated originals, how could any novel text that they generate be meaningful?
@@
In Part I, we defend bibliotechnism against this challenge. Using n-grams as a toy model, and working up to more complex modern LLMs, we show how even entirely novel text produced by LLMs may nevertheless derive its meaningfulness from the meaningfulness of inputs.
@@
We see this as a big step forward for bibliotechnism. But we go on (Part II) to present a new challenge for this proposal. Modern LLMs are not just capable of producing new sentences, they also seem to be able to use newly invented names apparently to refer to newly created objects. These new names cannot derive their reference from original text, since the name is not used in the data to refer to the relevant object. We argue that responding to this novel reference problem requires complicating bibliotechnism to such an extent that it calls into question the motivation for doing so.
@@
In particular, bibliotechnism offers a more complex, less predictive explanation of LLM behavior in the novel reference problem than can be given on the hypothesis that LLMs have beliefs, desires, and intentions. According to interpretationism in the philosophy of mind and cognitive science, a system has beliefs, desires and intentions if and only if its behavior is well explained by the hypothesis that it is rational and that it has such states (e.g., Dennett, 1971; Davidson, 1973, 1986; Dennett, 1989; McCarthy, 1979). If intrepretationism is true, then the novel reference problem provides evidence that LLMs have beliefs, desires, and intentions. More strongly, we argue that, for anyone who has some confidence in interpretationism, the novel reference problem provides evidence that LLMs do have these states.
2 Prior Work and Background
Prior Work
@@ 2 Prior Work and Background
A prominent line of argument has suggested that LLMs cannot produce reference without being “grounded” (e.g., Lake and Murphy, 2023; Bisk et al., 2020). Perhaps most influentially, Bender and Koller (2020) examine the question of whether LLMs understand language. They define “meaning” as a relation between expressions and communicative intents. They argue that LLMs cannot understand language in part because they cannot have perceptual contact with objects to which speakers intend to refer, and so cannot learn those speakers’ communicative intents. A natural inference from their discussion is that, owing to LLMs’ inability to understand language, they can also not produce meaningful text (cf. Titus, 2024).
@@
Mandelkern and Linzen (2023) observe important connections between this debate and semantic externalism, a view of meaning which has been dominant in the philosophy of language since the 1980s. On a standard view (Kripke, 1980; Putnam, 1975; Burge, 1986), people can refer to Shakespeare without having been directly in touch with Shakespeare, by belonging to a community whose overall use of this word stands in an appropriate causal relationship to the poet. Mandelkern and Linzen accordingly argue that whether LLMs can refer to Shakespeare comes down to whether LLMs “belong to our speech community” (cf. Ostertag, 2023a).
@@
The present paper goes beyond these earlier works by considering the relationship between the meaningfulness of LLM-generated text and the viability of bibliotechnism. Moreover, in contrast to earlier theoretical work, which does not clearly vindicate the claim that n-gram models produce meaningful text, we argue that n-gram models do produce such text. We build on this account to offer a new extension of bibliotechnism, which demonstrates that the view can accommodate the meaningfulness of even entirely novel text generated by LLMs. We then introduce the novel reference problem as a new challenge for bibliotechnism.
Background
@@
Three points about philosophical terminology will be important. Philosophers often distinguish between “reference” and “meaning”. As we use the terms, any expression that refers has a meaning, although many meaningful expressions do not refer. For simplicity, the only words that we take to refer are meaningful common and proper nouns. It is uncontroversial that a normal use of the word “Shakespeare” refers to Shakespeare (and is meaningful). By contrast, in our (stipulative) usage, the expression “was born” is typically meaningful, but does not refer.
Second, there is a difference between the word “Shakespeare” and particular inscriptions of this word. If the word “Shakespeare” is written on a blackboard five times, there are five inscriptions of this one word on the blackboard. We assume inscriptions of words can refer and be meaningful.
Third, and finally, we distinguish between “referring” that is done by an agent (“In his indirect way, Marlowe was referring to Queen Elizabeth.”), and referring that is done by particular inscriptions of words. Since our goal is to explore a view on which LLMs are not agents, we will be investigating the question of whether they can produce inscriptions which refer and are meaningful. We will not be assuming that they themselves can refer. [^1] Mandelkern and Linzen (2023) often move without comment between the question of whether a particular agent refers, and the question of whether particular inscriptions of words refer. But this distinction seems to us of key importance here, since it may take beliefs and desires to refer as an agent, but it does not take such attitudes to produce inscriptions which refer (as we will argue in a moment, photocopiers can do so, as can n-grams).
Part I: How to be a Bibliotechnist
3 Bibliotechnism and Derivative Meaning
@@
Gopnik and other bibliotechnists understand cultural technologies, like books and libraries, as tools for the transmission and dissemination of information, allowing the accumulation of knowledge over large stretches of space and time. These technologies are, crucially, not themselves responsible for new ideas or information.
@@
We will say that the original inscriptions, which were created by the author immediately, are instances of basic meaning and reference, while the other inscriptions are instances of derivative meaning and reference. We stipulate, as part of the definition of these terms, that it is only entities which have beliefs, desires, and intentions who produce inscriptions which refer or are meaningful basically. Beyond this stipulation, the distinction between basic and derivative reference and meaning is rough, but we will only deal with clear examples of each category in what follows.
According to bibliotechnism, LLMs do not have beliefs, desires, or intentions. So, according to bibliotechnism, LLMs can only produce inscriptions which refer or are meaningful derivatively.
@@
4 Causal History and Derivative Meaning
In this section, we argue that derivative reference and meaning can be achieved by an appropriate causal connection between PrimaryData and GeneratedText, and show how this vindicates the idea that n-grams can produce derivatively meaningful inscriptions of individual words.
@@ 4 Causal History and Derivative Meaning
This observation already shows that large-n n-gram models—which sample from a distribution conditional on the previous $n-1$ words—can produce meaningful inscriptions. For sufficiently large $n$ (e.g. 1000), such models simply copy particular inscriptions from their PrimaryData. So, like a photocopier, they produce meaningful output.
Matters are less straightforward for unigram models, which sample from a distribution of single words. We think of such models as implemented by taking all of their PrimaryData, choosing word-inscriptions from the PrimaryData at random, and then copying the chosen inscription. The inscriptions of individual words the n-gram then produces are again just like those of a copier (or of the large-n model): they have a direct causal connection to the original inscriptions. As a result, if the model produces an inscription of “Shakespeare”, this inscription will be meaningful (and refer) derivatively, piggybacking on the meaning and reference of the original inscription of this word.
In this case, however, there is a new phenomenon, not exhibited in the case of the photocopier or large-n n-gram. Each of the inscriptions of the individual words produced by the unigram will be meaningful, but it does not seem that inscriptions of complex expressions formed from these words will be. The vast majority of the time, the string of words the model produces will be gibberish, and uncontroversially meaningless. At low odds the unigram model will produce a “reasonable” string like “Shakespeare was born in 1564”. With even lower odds, such a reasonable string will be produced by copying an original token string. [^2] Recall that we take the model to be implemented by randomizing over inscriptions and copying the particular inscription it draws, so copying a particular string of inscriptions is a different event than producing an inscription that happens already to be in the data. But even in these latter cases, the fact that a string which could be meaningful is produced is a fluke, a complete accident. As a result, we judge that, even when the model produces an inscription of a string that would be meaningful if produced by a human in a normal way, the inscription the model produces is not meaningful. The generated text is not appropriately causally connected to the PrimaryData to inherit the “glue” that binds the words in the complex expression together. One way to put this point would be to say that, while the resulting inscription may look like an inscription of a sentence (and a human who sees it may be able to conjure up a meaning associated with it), it is not really an inscription of a sentence, but just an inscription of words which could (in different circumstances) have made up a sentence. These inscriptions are like sand dunes blown into the shape of a sentence—or like Mandelkern and Linzen’s “ants formicating meaninglessly in the sand”.
5 Derivatively Meaningful Novel Expressions
@@
We will argue that it can, by arguing that inscriptions of complex expressions can be derivatively meaningful even if two distinct causal pathways are involved in their production: a first (discussed above), which connects new inscriptions of individual words to originals in the data; and a second (new to this section), which guarantees that outputs possess higher-level features of expressions in the original PrimaryData to serve as a kind of “glue” binding inscriptions of meaningful words.
@@
To see the basic idea, consider a rudimentary model, which, when fed a sentence, finds any inscriptions of names in the sentence (searching on the basis of a database) and then replaces each of these names uniformly with a name drawn at random from the distribution of all names in its PrimaryData. This model plausibly produces not just inscriptions of individual words which are derivatively meaningful, but an inscription of a new sentence which would, as a whole, be derivatively meaningful. For instance, if we gave this model our Shakespeare sentence, and it produced “Barack Obama was born in 1564” this inscription would be false, but meaningful, even if it has never been contemplated before by a human. Here, the causal history of the context and the causal history of the individual name printed are different, but the whole expression would still be derivatively meaningful. Plausibly, this is because the operation as a whole is causally sensitive to a structural, higher-level features of its input, and in particular, that it reliably produces an output sentence that preserves the grammatical structure of its input sentence. A rough test for causal sensitivity in this sense (though not a necessary or sufficient condition) appeals to counterfactual sensitivity: (i) if the model were given PrimaryData exhibiting property $P$, would it reliably produce outputs which exhibit property $P$?; and (ii) if the model were given PrimaryData which does not exhibit property $P$, would it reliably produce outputs which do not exhibit property $P$? The rudimentary model just described passes this test, and is in any case intuitively causally sensitive to the structure of its input sentence. It is in part owing to this causal sensitivity that the model can produce inscriptions of novel sentences which are derivatively meaningful, since the output sentences inherit not only the meaningfulness of their constituents, but also their form from the PrimaryData.
This discussion provides one example where novel text can nevertheless be derivatively meaningful. It also suggests that, in general, if a model is causally sensitive to relevant high-level features of its PrimaryData, in such a way as to transmit those features to its GeneratedText, it is possible that even entirely novel GeneratedText can be derivatively meaningful. The question then becomes: is there such a high-level feature in the case of modern LLMs?
One proposed answer might focus on grammaticality. The rudimentary model just described preserves the grammaticality of its input. Accordingly, one might think that it would suffice for an LLM to produce meaningful output, if it is causally sensitive to the grammaticality of its PrimaryData in such a way that it reliably produces grammatical GeneratedText. But this proposal is not correct in general, because not all grammatical sentences are meaningful. So, even if a process reliably produces grammatical sentences with derivatively meaningful sub-expressions, the whole sentence could fail to be derivatively meaningful, if the resulting sentence were grammatical but not meaningful. In fact, as Gulordava et al. (2018) show, even the results of a simple operation which replaces meaningful constituents with other meaningful constituents can often lead to sentences that are grammatical but meaningless, e.g.: “You apply the toy and serve fighter hair into the blackmail .” If LLM-produced text is meaningful, it cannot be only in virtue of causal sensitivity to the grammaticality of its input.
These observations lead us to suggest that the relevant high-level property is not grammaticality but intelligibility. As we will understand this notion, the intelligibility of an expression does require that it be at least quasi-grammatical (sentences with minor grammatical errors are often perfectly intelligible). But, as we have seen, even perfect grammaticality does not on its own suffice for intelligibility. We will not offer further analysis of the notion of intelligibility, and the boundary between intelligility and unintelligibility may be vague, but there are clear cases on both sides: “Shakespeare was born in 1564” is intelligible, while “You apply the toy and serve fighter hair into the blackmail ” is not. We suggest that, if LLMs are causally sensitive to the intelligibility of their input, in such a way as to produce intelligible outputs when given intelligible inputs, then intelligible complex expressions in their GeneratedText will be derivatively meaningful, with individual words inheriting the meanings of the original inscriptions from which they were “copied”, and whole expressions inheriting the “glue” of intelligibility from the PrimaryData.  Earlier we said that counterfactual sensitivity is a rough test for causal sensitivity. This test gives us evidence that modern LLMs are in fact appropriately causally sensitive to the intelligibility of their PrimaryData. First, these models overwhelmingly produce text which is clearly intelligible. Second, it seems extremely plausible that if LLMs were trained on gibberish, they would output gibberish. These two claims at least point toward the verdict that they are causally sensitive in the requisite sense.
@@
Intelligibility in our sense does not require truth or even sufficiently reliable production of the truth. False sentences like “Shakespeare was born in 2023” are perfectly intelligible. Even the best LLMs at the time of writing are known to confabulate or fabricate information. But getting a fact wrong (e.g., saying Shakespeare was born in 2023 instead of 1564) is importantly different than producing incoherent and in particular unintelligible responses. If an LLM reliably responded to queries about Shakespeare’s birth with gibberish, this would at least be some evidence that it is not in fact causally sensitive to the intelligibility of its data in such a way as to generate derivatively meaningful complex expressions.
It is instructive to compare LLM-generated novel text to text generated by bigram or trigram models. Unlike unigrams, bigram and trigram models fairly reliably copy short complex phrases. Indeed, such models might be statistically likely to combine expressions in ways that might seem meaningful as a whole, because they are causally sensitive to certain features of the patterns of combination of these words in the data. For instance, a bigram model that outputs “Shakespeare wrote plays” does so in part because it is sensitive to the fact that “wrote” is a likely continuation for “Shakespeare” and that “plays” is a likely continuation for ”wrote”. But, the only causal connection between its production of “Shakespeare” and its production of “plays” is mediated by the verb “wrote”. Given this fact, we judge that the causal story about its production of this sentence does not preserve an appropriate connection between all of the parts of the sentence. Accordingly, even when the model produces strings of sentence-length that are grammatical, and even when individual phrases may be judged meaningful, it seems that, as with unigram models, longer sentences should probably not be understood as meaningful (although it is much more of a borderline case). These sentences lack the straightforward “copy property” of higher-n n-gram models but also are not produced in a way that is causally sensitive to relevant structural features (and in particular intelligibility), as modern LLMs seem to be.
@@
Before closing this discussion, we want to offer one important clarification about the basis of our judgment that unigrams do not produce derivatively meaningful complex expressions. The basis for this judgment is not the fact that n-grams are only trained on individual words. It is instead because the structure and training of n-grams does not lead to causal sensitivity to relevant high-level features of their PrimaryData. To put this another way: we are not interested in a narrow form of “input-sensitivity”, but instead in a broader notion of causal sensitivity, partly captured by the test of counterfactual sensitivity.
@@
This contrast can be illustrated by considering again a photocopier. The fact that a photocopier responds to (say) one or another aspect of the ink used to write original letters is irrelevant to the question of whether the inscriptions it produces are meaningful. As long as its underlying low-level mechanism leads to causal sensitivity to the right high-level features—as evidenced in this case by the fact that it reliably produces inscriptions of words when it is fed words, and reliably does not produce words when it is not fed words—the inscriptions it produces will be derivatively meaningful.